Project

General

Profile

NDOS Commissioning Meeting 2011/09/12

Attendance: Jon, Martin, Mat, Gavin, Enrique, Mark Mathis, Denis, Kurt, Andrew, Tom, Rafael, Alex, Minerba, Susan, Jan, Leon,

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • Jon spoke with Sam Childress re: NuMI target. They are on-schedule with the target, and plan to shutdown on Thurs. morning to replace the target. Should take ~8 days to get back to normal operations.
  • Smooth operations over the weekend.
  • Mat:
    • took another DSO scan on Friday and found 5 more cooled APDs that are masked out now due to noise (1-2, 1-3, top of diblock 3). Of the golden channels that we 35% golden are still working, others: 20% still working. Have seen several APDs go bad after several months.
    • Jon (to Leon): is there still more to learn from cooling? Leon: yes, we'd like to know the TECC failure rate.

DAQ Status & Plans

  • Leon:
    • trip to Japan:
      • good news: much better understanding of their procedures.
      • bad news: will get ~250 APDs (all coated), but limiting factor is wafer production which is driving lead time. Starting wafers now, but it will be 3.5 months before 1st large batch is ready. Should get another 10 "mechanical" prototypes that will be silicone coated in ~6 weeks (end of Oct.) that would be put on detector after testing at CalTech.
    • seal-tests have been completed on the 20 APDs at CalTech and all were fine (not shipped out yet). Has not heard back from Stuart Mufson about sealing of heatsinks.
  • Andrew: trying to work out interface between dispatcher and ART framework. Trying to get new event display up and running. Test software is being built on datamon. Tests will begin tomorrow morning.
  • Kurt: put together GUI bit map mask for Cfg editor

DCS Status & Plans

  • Martin:
    • Gennadiy developed on a new GUI interface which is now being rolled-out
    • Athans would like to give a short presentation at the next commissioning meeting
    • Martin will work with Mat on maintaining cooling bit masks.

AOB

  • Assessment of plan to rearrange electronics on detector:
    • Mat: how long would it take to rearrange? Needs some time to think about this.
    • Leon: idea of coating existing APDs with paralene (sp?). Will talk to John Cooper about this.
    • Tricia: probably don't want to completely sacrifice muon tracking for this.
    • Minerba (presentation):
      • studied effect of having only fid. volume instrumented
      • number of NC events found ~doubles
      • number of CC nu_mu events goes down slightly (although number of QE events goes up slightly, both RES and DIS events drop).
      • track length information is lost, but total PE near the vertex goes up